Circular No. 7059

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998es IN NGC 632 K. Ayani, Bisei Astronomical Observatory (BAO); and H. Yamaoka,

Kyushu University, report: “A spectrum of SN 1998es (cf. IAUC 7050, 7054) was obtained with BAO 1.01-m telescope on Nov. 19.6 UT. Compared with the spectrum reported on IAUC 7054 (see http://cfa-www.

harvard.edu/cfa/oir/Research/supernova/spectra/sn98es.gif), a Si

II absorption feature around 612 nm (on the rest frame of the host galaxy) has deepened but is still weaker than that of typical type-Ia supernovae. This spectral evolution is very similar to the intrinsically bright type-Ia supernova 1991T, and our spectrum shows that SN 1998es would be around maximum on the observed time. Narrow Na I D absorption lines indicate that the recession velocity of the host galaxy is ∼ 3200 km/s, and that SN 1998es is slightly reddened. The expansion velocity derived from the shift of Si II absorption minimum is ∼ 11 000 km/s, which is smaller than reported by Jha et al. (IAUC 7054). Our spectrum is locat-

ed at ftp://ftp.kusastro.kyoto-u.ac.jp/pub/vsnet/SNe/sn1998es/
sn98es550.gif.”

COMET C/1998 U5 (LINEAR)

Ephemeris extension to IAUC 7045 (elements MPC 32866):

1998TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Nov. 23 23^{h}22.^{m}14 +51^{o}53.^{′}5 0.507 1.313 119.^{o}4 40.^{o}9 7.7
25 22 52.34 +47 44.8 0.544 1.303 113.5 44.0 7.8
27 22 30.67 +43 50.4 0.586 1.294 108.0 46.5 8.0
29 22 14.56 +40 19.4 0.633 1.285 102.9 48.4 8.1
Dec. 1 22 02.31 +37 13.3 0.683 1.278  98.2 49.8 8.2
3 21 52.79 +34 31.0 0.736 1.270  94.0 50.7 8.4
5 21 45.24 +32 09.7 0.790 1.264  90.1 51.2 8.5
7 21 39.17 +30 06.7 0.846 1.258  86.4 51.4 8.6
9 21 34.21 +28 19.2 0.902 1.253  83.1 51.3 8.8
11 21 30.12 +26 45.0 0.959 1.249  79.9 50.9 8.9
13 21 26.70 +25 22.1 1.016 1.245  77.0 50.4 9.0

Visual m_{1} and coma-diameter estimates (B = binoculars): Nov. 16.01 UT, 7.6, 14′ (J. Bortle, Stormville, NY, 10×50 B); 16.91, 8.1, 10′ (K. Hornoch, Lelekovice, Czech Republic, 10×80 B); 18.72, 8.0, 7′.6 (W. Hasubick, Buchloe, Germany, 25×100 B); 20.16, 7.8, 15′ (R. Keen, Mt. Thorodin, CO, 7×50 B); 23.92, 7.8, 5′ (A. Baransky, Kiev, Ukraine, 12×80 B).

1998 November 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green